# ViHateT5: Enhancing Hate Speech Detection in Vietnamese With a Unified Text-to-Text Transformer Model

### v1

@@ Abstract
Recent advancements in hate speech detection (HSD) in Vietnamese have made significant progress, primarily attributed to the emergence of transformer-based pre-trained language models, particularly those built on the BERT architecture. However, the necessity for specialized fine-tuned models has resulted in the complexity and fragmentation of developing a multitasking HSD system. Moreover, most current methodologies focus on fine-tuning general pre-trained models, primarily trained on formal textual datasets like Wikipedia, which may not accurately capture human behavior on online platforms. In this research, we introduce ViHateT5, a T5-based model pre-trained on our proposed large-scale domain-specific dataset named VOZ-HSD. By harnessing the power of a text-to-text architecture, ViHateT5 can tackle multiple tasks using a unified model and achieve state-of-the-art performance across all standard HSD benchmarks in Vietnamese. Our experiments also underscore the significance of label distribution in pre-training data on model efficacy. We provide our experimental materials for research purposes, including the VOZ-HSD dataset [^1] https://huggingface.co/datasets/tarudesu/VOZ-HSD, pre-trained checkpoint [^2] https://huggingface.co/tarudesu/ViHateT5-base, the unified HSD-multitask ViHateT5 model [^3] https://huggingface.co/tarudesu/ViHateT5-base-HSD, and related source code on GitHub [^4] https://github.com/tarudesu/ViHateT5 publicly.
@@ 1 Introduction
Different languages have unique forms of harmful expressions, necessitating specific text-processing methodologies for developing HSD systems. In the context of English, one of the most prevalent languages, there exist several effective strategies for addressing HSD-related tasks, such as employing machine learning models Abro et al. (2020) or deep learning models for identifying harmful content Badjatiya et al. (2017); Zimmerman et al. (2018). Furthermore, transfer learning approaches have garnered considerable research interest, showcasing remarkable performance in hate speech detection tasks Ali et al. (2022); Mozafari et al. (2020). In the case of low-resource languages, numerous studies have been conducted, yielding promising results in tackling this issue Bigoulaeva et al. (2021); Nkemelu et al. (2022); Arango Monnar et al. (2022).
@@
Vietnamese, considered a low-resource language, has seen limited research in natural language processing concerning large-scale datasets and pre-trained models. Recent efforts in hate speech detection tasks based on Vietnamese language characteristics have yielded significant achievements Vu et al. (2020); Luu et al. (2021); Hoang et al. (2023). However, current state-of-the-art models only fine-tune general transformer-based models, which may have been pre-trained on formal textual data sources Nguyen et al. (2020). Moreover, even pre-trained on social media text, models like ViSoBERT Nguyen et al. (2023) still necessitate separate fine-tuning for specific tasks, resulting in system fragmentation.
In this paper, we present a new approach to address the existing limitations of HSD systems in Vietnamese. The contributions of this research are outlined as follows:
  • A novel domain-specific model named ViHateT5 is presented in this study to address HSD-related problems in the Vietnamese language. Unlike prior endeavors, which merely fine-tune general transformer models, our innovative model was explicitly trained on a specific dataset derived from social media texts called VOZ-HSD with 10M+ comments with generated labels.
  • A unified T5-based model by fine-tuning the pre-trained ViHateT5 model advances the state-of-the-art performance on all available HSD benchmark datasets in Vietnamese.
@@
This paper is organized into distinct sections. Section 2 examines relevant prior research on hate speech detection tasks in Vietnamese. Subsequently, Section 3 introduces ViHateT5, our text-to-text model, covering its automatically generated pre-training dataset, pre-training methodologies, and fine-tuning for downstream tasks. Section 4 presents experimental results obtained by comparing various baseline methods with our proposed ViHateT5 model across a range of hate speech detection-related tasks and address discussions. Section 5 concludes the paper with a summary of our findings. Section 6 addresses the current limitations of our proposed method, while Section 7 provides ethical statements related to our research.
@@ 2 Related Work
On the other hand, despite the remarkable performance achieved by current BERT-based approaches, the necessity persists to develop separate systems tailored to individual tasks. Addressing this issue of fragmentation, a unified text-to-text-based model on T5 architecture Raffel et al. (2023) such as FT5 for English or mFT5 Ranasinghe and Zampieri (2023) has demonstrated its effectiveness in amalgamating multiple tasks ranging from syllable-level to sentence-level-based HSD challenges. Hence, in this study, we introduce ViHateT5, a pre-trained text-to-text model designed explicitly for the hate speech domain, aiming to streamline complex separate systems while ensuring optimal performance in addressing HSD issues in Vietnamese.
@@ 3.1 Automated Pre-training Data Creation
Vietnamese is a low-resource language, which results in a need for more extensive datasets for training targeted language models, particularly in specific NLP tasks. In this research, we present a significant Vietnamese hate speech classification dataset alongside an automated data annotation system. Figure 2 illustrates the entire process, which includes several modules.
@@
Automated Data Labeling. Utilizing the selected HSD Classifier, we automatically label all textual data within the raw dataset. The resultant dataset comprises approximately 10 million user comments annotated with hate speech labels. According to the statistics, there are over 500K+ comments labeled as containing harmful content, constituting a significant portion of the total dataset. Notably, the dataset still maintains a substantial number of comments labeled as hate speech, particularly when compared to label distributions observed in the previous study by Luu et al. (2021) focusing on the Vietnamese hate speech detection task. We designate the final dataset as VOZ-HSD, indicating its purpose for hate speech detection and its origin from VOZ.
@@ 3.2 Model Pre-training
Inspired by the success of domain-specific pre-training and the efficacy of text-to-text models in addressing HSD tasks Ranasinghe and Zampieri (2023), we embark on the pre-training of the ViHateT5 model, leveraging the T5 architecture. The constructed VOZ-HSD dataset is employed as the pre-training dataset, comprising samples extracted from real-life comments.
@@ 3.3 Model Fine-tuning
Toxic Speech Detection (ViCTSD). The ViCTSD dataset Nguyen et al. (2021) was initially formulated to identify constructiveness and toxicity in user comments. However, this study focuses solely on detecting toxicity within the comments. Originating from online news websites, where users are typically middle-aged individuals who often express themselves in formal styles, the level of offensiveness may not be as overt as in other datasets. Consequently, conducting toxicity detection on this dataset challenges language models.
@@ 4.1 Data
This section outlines the experiments conducted, including details regarding the training data utilized, model configurations, baselines, evaluation metrics, and the ensuing results. We present the performance of our proposed ViHateT5 model across multiple HSD tasks in Vietnamese, providing insights into its efficacy relative to other previous state-of-the-art approaches. Note that no specific data pre-processing is applied to any models to ensure fairness.
Pre-training Data. We use raw texts from the VOZ-HSD dataset to pre-train the ViHateT5 models. As labeled by the HSD Classifier, raw texts in the VOZ-HSD dataset with generated labels can be helpful for data analysis and experiments with different proportions of hate-labeled samples.
@@ 4.2 Model Setup
We follow the original pre-training strategy outlined for the T5 model Raffel et al. (2023) to pre-train our ViHateT5. Both training and validation are conducted with a batch size of 128. The pre-training process is executed over 20 epochs, employing the Adam optimizer with a lower learning rate set at 5e-3. Additionally, a weight decay of 0.001 is applied, with the initial 2,000 steps designated for warm-up during training.
@@ 4.3.1 Multilingual Pre-trained Models
BERT Devlin et al. (2019) represents a landmark achievement in NLP tasks. Unlike previous models, BERT leverages the Transformer architecture for pre-training on extensive corpora, simultaneously leveraging both left and right context in every layer. This bidirectional representation grants BERT a contextual understanding of language surpassing prior methods. Consequently, BERT obtained state-of-the-art performance on various NLP tasks. Furthermore, BERT requires minimal fine-tuning for downstream tasks, demonstrating its remarkable generalizability. As a result, BERT has become a fundamental component in NLP pipelines. Besides the original English-only pre-trained model, the multilingual version supporting 100+ languages, including Vietnamese, has been released to solve problems in various languages.
DistilBERT, introduced by Sanh et al. (2019), is a compact Transformer model derived from the popular BERT architecture. Knowledge distillation effectively captures the knowledge of a larger pre-trained model while drastically reducing its size and computational complexity. Despite its smaller size, DistilBERT exhibits remarkable performance, retaining over 95% of BERT’s accuracy on the GLUE language understanding benchmark. The multilingual version of DistilBERT has also been publicly released.
XLM-RoBERTa developed by Conneau and Lample (2019), stands as a prominent achievement in multilingual language modeling. Leveraging a Transformer architecture and trained on a massive dataset of 2.5TB filtered CommonCrawl text across 100 languages, it employs a masked language modeling (MLM) objective to learn cross-lingual representations. This approach surpasses previous models like mBERT in diverse tasks. As a result, XLM-R offers a compelling solution for cross-lingual tasks, pushing the boundaries of multilingual language understanding.
@@ 4.3.2 Monolingual Pre-trained Models
PhoBERT, known as one of the first public large-scale monolingual language models for Vietnamese, has achieved excellent performance on various NLP tasks Nguyen et al. (2020) since its emergence. PhoBERT initially comes in two sizes, base and large, catering to different needs and computing resources. Built upon the successful RoBERTa architecture, PhoBERT offers improved robustness and performance. PhoBERT is now available in three versions, including base, large, and base version-2nd.
@@ 4.6 Discussion
The effectiveness of pre-training a transformer model on a domain-specific dataset was further validated by the ViSoBERT model, which demonstrated superior performance across various social media benchmark datasets Nguyen et al. (2023).
How Pre-training Data Affects ViHateT5. We also assess how varying the data ratio in pre-training data affects our proposed models. This evaluation involves pre-training under different data conditions: utilizing full-data samples as in this study, employing a balanced-label pre-trained model with equal samples for both labels and utilizing a hate-only pre-trained model where only hate labels are retained for pre-training.
@@ B.1 Model Pre-training
Table 7: Model settings for pre-training ViHateT5 variants. Note that all pre-trained models were trained on a low-resource setting with a single GPU NVIDIA A6000.
@@ Appendix C What is inside the VOZ-HSD dataset?
Table 10 illustrates the distribution of topic text data within the VOZ-HSD dataset. It is evident that we have gathered a wide range of conversation topics, indicating that the dataset is not skewed towards any particular domain and closely reflects real-life textual content. Previous studies by Nguyen et al. Nguyen et al. (2023) have further demonstrated that even with a limited dataset size of only 1GB in an uncompressed format for pre-training a transformer on social media texts, the model can still exhibit strong performance across multiple tasks, achieving state-of-the-art results.

### v2

@@ Abstract
Recent advancements in hate speech detection (HSD) in Vietnamese have made significant progress, primarily attributed to the emergence of transformer-based pre-trained language models, particularly those built on the BERT architecture. However, the necessity for specialized fine-tuned models has resulted in the complexity and fragmentation of developing a multitasking HSD system. Moreover, most current methodologies focus on fine-tuning general pre-trained models, primarily trained on formal textual datasets like Wikipedia, which may not accurately capture human behavior on online platforms. In this research, we introduce ViHateT5, a T5-based model pre-trained on our proposed large-scale domain-specific dataset named VOZ-HSD. By harnessing the power of a text-to-text architecture, ViHateT5 can tackle multiple tasks using a unified model and obtain state-of-the-art performance on all benchmark HSD datasets in Vietnamese. The experiments also underscore the significance of label distribution in pre-training data on model efficacy. We provide our experimental materials for research purposes, including the VOZ-HSD dataset [^1] https://huggingface.co/datasets/tarudesu/VOZ-HSD, pre-trained checkpoint [^2] https://huggingface.co/tarudesu/ViHateT5-base, the unified HSD-multitask ViHateT5 model [^3] https://huggingface.co/tarudesu/ViHateT5-base-HSD, and related source code on GitHub [^4] https://github.com/tarudesu/ViHateT5 publicly.
@@ 1 Introduction
Different languages have unique forms of harmful expressions, necessitating specific text-processing methodologies for developing HSD systems. In the context of English, one of the most prevalent languages, there exist several effective strategies for addressing HSD-related tasks, such as employing machine learning models Abro et al. (2020) or deep learning models for identifying harmful content Badjatiya et al. (2017); Zimmerman et al. (2018). Furthermore, transfer learning approaches have garnered considerable research interest, showcasing the effectiveness in solving hate speech detection tasks Ali et al. (2022); Mozafari et al. (2020). In the case of low-resource languages, numerous studies have been conducted, yielding promising results in tackling this issue Bigoulaeva et al. (2021); Nkemelu et al. (2022); Arango Monnar et al. (2022).
@@
Vietnamese, considered a low-resource language, has seen limited research in the field of NLP concerning high-quality datasets and pre-trained models. Recent efforts in hate speech detection tasks based on Vietnamese language characteristics have yielded significant achievements Vu et al. (2020); Luu et al. (2021); Hoang et al. (2023). However, current state-of-the-art models only fine-tune general transformer-based models, which may have been pre-trained on formal textual data sources Nguyen et al. (2020). Moreover, even pre-trained on social media text, models like ViSoBERT Nguyen et al. (2023) still necessitate separate fine-tuning for specific tasks, resulting in system fragmentation.
This paper presents a novel approach to address the existing limitations of HSD systems in Vietnamese. The contributions of this research are outlined as follows:
  • A domain-specific model named ViHateT5 is presented in this study to address HSD-related problems in the Vietnamese language. Our innovative model was explicitly trained on a specific dataset derived from social media texts called VOZ-HSD with 10M+ comments with generated labels, designed to perform multiple tasks in a unified framework.
  • A unified T5-based model, obtained by fine-tuning the pre-trained ViHateT5 model, advances the state-of-the-art performance on all available HSD benchmark datasets in Vietnamese.
@@
This paper is structured into distinct sections. Section 2 examines relevant prior research on Vietnamese HSD-related tasks. Subsequently, Section 3 introduces ViHateT5, our text-to-text model, covering its automatically generated pre-training dataset, pre-training methodologies, and fine-tuning for downstream tasks. Section 4 presents experimental results obtained by comparing various baseline methods with our proposed ViHateT5 model across a range of hate speech detection-related tasks and address discussions. Section 5 concludes the paper with a summary of our findings. Section 6 addresses the current limitations of our proposed method, while Section 7 provides ethical statements related to our research.
@@ 2 Related Work
On the contrary, despite the remarkable performance achieved by current BERT-based approaches, developing separate systems tailored to individual tasks is necessary. Addressing this issue of fragmentation, a unified text-to-text-based model on T5 architecture Raffel et al. (2023) such as FT5 for English or mFT5 Ranasinghe and Zampieri (2023) has demonstrated its effectiveness in amalgamating multiple tasks ranging from syllable-level to sentence-level-based HSD challenges. Hence, in this study, we introduce ViHateT5, a pre-trained text-to-text model designed explicitly for the hate speech domain, aiming to streamline complex separate systems while ensuring optimal performance in addressing HSD issues in Vietnamese.
@@ 3.1 Automated Pre-training Data Creation
Vietnamese is a low-resource language, which results in a need for more extensive datasets for training targeted language models, particularly in specific NLP tasks. In this research, we present a massive Vietnamese hate speech classification dataset alongside an automated data annotation system. Figure 2 illustrates the entire process, which includes several modules.
@@
Automated Data Labeling. Utilizing the selected HSD Classifier, we automatically label all textual data within the raw dataset. The resultant dataset comprises approximately 10 million user comments annotated with hate speech labels. According to the statistics, there are over 500K+ comments labeled as containing the harmful content, constituting a significant portion of the total dataset. Notably, the dataset still maintains a substantial number of comments labeled as hate speech, particularly when compared to label distributions observed in the previous study by Luu et al. (2021) focusing on the Vietnamese hate speech detection task. We designate the final dataset as VOZ-HSD, indicating its purpose for hate speech detection and its origin from VOZ.
@@ 3.2 Model Pre-training
Domain-specific pre-training and the efficacy of text-to-text models in addressing HSD tasks Ranasinghe and Zampieri (2023) have attained significant achievements. Following that, we embark on the pre-training of the ViHateT5 model, leveraging the T5 architecture. The constructed VOZ-HSD dataset is employed as the pre-training dataset, comprising samples extracted from real-life comments.
@@ 3.3 Model Fine-tuning
Toxic Speech Detection (ViCTSD). The ViCTSD dataset Nguyen et al. (2021) was initially formulated to identify constructiveness and toxicity in user comments. Originating from online news websites, where users are typically middle-aged individuals who often express themselves in formal styles, the level of offensiveness may not be as overt as in other datasets. Consequently, conducting toxicity detection on this dataset challenges language models.
@@ 4.1 Data
In this section, we outline the experiments conducted, including details regarding the training data utilized, model configurations, baselines, evaluation metrics, and the ensuing results. We illustrate our proposed ViHateT5 model performance across multiple HSD tasks in Vietnamese, providing insights into its efficacy relative to other previous state-of-the-art approaches. Note that no specific data pre-processing is applied to any models to ensure fairness.
Pre-training Data. The raw texts from the VOZ-HSD dataset are used for the pre-training phase of the ViHateT5 model. As labeled by the HSD Classifier, raw texts in the VOZ-HSD dataset with generated labels can be helpful for data analysis and experiments with different proportions of hate-labeled samples.
@@ 4.2 Model Setup
We follow the original pre-training strategy outlined for the T5 model Raffel et al. (2023) to pre-train our ViHateT5. Both training and validation are conducted with 128 batch size. The process for pre-training is executed over 20 epochs, employing the Adam optimizer with a lower learning rate set at 5e-3. Additionally, a weight decay of 0.001 is applied, with the initial 2,000 steps designated for warm-up during training.
@@ 4.3.1 Multilingual Pre-trained Models
BERT. Devlin et al. (2019) introduced their remarkable pre-training language model called BERT, representing a landmark achievement in NLP tasks. Unlike previous models, BERT leverages the Transformer architecture for pre-training on extensive corpora, simultaneously leveraging both left and right context in every layer. This bidirectional representation grants BERT a contextual understanding of language surpassing prior methods. Consequently, BERT performed excellently on various NLP tasks. Furthermore, BERT only needs minimal fine-tuning for downstream tasks, demonstrating its remarkable generalizability. As a result, BERT has become a fundamental component in NLP pipelines. Besides the original English-only pre-trained model, the multilingual version supporting 100+ languages, including Vietnamese, has been released to solve problems in various languages.
DistilBERT, introduced by Sanh et al. (2019), is a compact pre-trained language model derived from the popular BERT architecture. Knowledge distillation effectively captures the knowledge of a larger pre-trained model while drastically reducing its size and computational complexity. Despite its smaller size, DistilBERT exhibits remarkable performance, retaining over 95% of BERT’s accuracy on the GLUE language understanding benchmark. The multilingual version of DistilBERT has also been publicly released.
XLM-RoBERTa, developed by Conneau and Lample (2019), stands as a prominent achievement in multilingual language modeling. Leveraging a Transformer architecture and trained on a massive dataset of 2.5TB filtered CommonCrawl text across 100 languages, it employs a masked language modeling (MLM) objective to learn cross-lingual representations. This approach surpasses previous models like mBERT in diverse tasks. As a result, XLM-R offers a compelling solution for cross-lingual tasks, pushing the boundaries of multilingual language understanding.
@@ 4.3.2 Monolingual Pre-trained Models
PhoBERT, known as one of the public large-scale language models for Vietnamese, has achieved excellent performance on various NLP tasks Nguyen et al. (2020) since its emergence. PhoBERT initially comes in two sizes, base and large, catering to different needs and computing resources. Built upon the successful RoBERTa architecture, PhoBERT offers improved robustness and performance. PhoBERT is now available in three versions, including base, large, and base version-2nd.
@@ 4.6 Discussion
How Pre-training Data Affects ViHateT5. The effectiveness of pre-training a transformer model on a domain-specific dataset was further validated by the ViSoBERT model, which demonstrated superior performance across various social media benchmark datasets Nguyen et al. (2023). In this work, we assess how varying the data ratio in pre-training data affects our proposed models. This evaluation involves pre-training under different data conditions: utilizing full-data samples as in this study, employing a balanced-label pre-trained model with equal samples for both labels and utilizing a hate-only pre-trained model where only hate labels are retained for pre-training.
@@ B.1 Model Pre-training
Table 7: Model settings for pre-training ViHateT5 variants. Note that all pre-trained models were trained on a limited-resource setting with a single GPU NVIDIA A6000.
@@ Appendix C What is inside the VOZ-HSD dataset?
Table 10 illustrates the distribution of topic text data within the VOZ-HSD dataset. It is evident that we have gathered a wide range of conversation topics, indicating that the dataset is not skewed towards any particular domain and closely reflects real-life textual content. Previous studies by Nguyen et al. (2023) have further demonstrated that even with a limited dataset size of only 1GB in an uncompressed format for pre-training a transformer on social media texts, the model can still exhibit strong performance across multiple tasks, achieving state-of-the-art results.